OpenStudy (anonymous):

2 questions!! Variables on both sides. Solve each equation. 1) n-4 = -11+ 2n + 5 -3 2) 1+ a = 2a + 6 -2a

OpenStudy (stefrheart):

1) n-4=-9+2n -n-4=-9 -n=-5 n=5